OpenStudy (agent_a):

Use the method of cylindrical shells to find the volume generated by rotating the region bounded by the given curves about the y-axis. y=x^2 , y=0 , x=1 , x=2

OpenStudy (amistre64):

if we are doing shells we will need to know the formula for it

OpenStudy (agent_a):

This is the equation that I used: \[\int\limits_{1}^{2}2\pi(2-x)(x^2) dx\] ...based on the formula: \[\int\limits_{a}^{b}2\pi xf(x)dx \]

OpenStudy (agent_a):

I solved the problem already. I just wanted to know if my answer is correct or not.

OpenStudy (agent_a):

Cylindrical shells have a standard formula, depending on the axis of rotation. If the rotation is done about the y-axis or x=*something* , the radius is in terms of x (vertical cylinder): \[\int\limits_{a}^{b}2\pi xf(x)dx\] If the rotation is done about the x-axis or y=*something*, the radius is in terms of y (horizontal cylinder): \[\int\limits_{c}^{d}2\pi yg(y)dy\]

zepdrix (zepdrix):

Hmm the radius of your shell looks strange.. \(\large (2-x)\). Wouldn't it simply be \(\large x\) since we're spinning from the y-axis?

OpenStudy (amistre64):

i tend to go about this as:\[2\pi ~rh\]is the area of one of the shells and then define h and r in their proper terms

OpenStudy (amistre64):

y=x^2 , y=0 , x=1 , x=2